How many gallons of a 20% antifreeze solution and a 10% antifreeze solution must be mixed to obtain 40 gallons of a 16% antifreeze solution?
Answer by checkley77(12844)

.20x+.10(40-x)=.16*40
.20x+4-.10x=6.4
.10x=6.4-4
.10x=2.4
x=2.4/.10
x=24 gallons of 20% antifreeze is used.
440-24=16 gallons of 10% antifreeze is used.
Proof:
.20*24+.10(40-24)=.16*40
4.8+.10*16=6.4
4.8+1.6=6.4
6.4=6.4